a bag contains 160 coins, there are coins of 10 and 25 cents. find the number of coin of 25 cents, if there is $26.80 in total
Answer by fractalier(6550)

Call the number of dimes and quarters, x and y. Then we can write
x + y = 160 and the value equation is
10x + 25y = 2680 (we work in cents)
Now multiply the top equation by ten and subtract it from the bottom...we get
10x + 25y = 2680
-(10x + 10y = 1600)
-------------------
15y = 1080
and dividing by 15...
y = 72 quarters...that means there are
x = 160 - 72 = 88 dimes